# What to do if language models disagree? Black-box model ensembling for textual and visual question answering

### v1

@@ 4.6 Ablation Studies
Is InfoSel robust to the base models’ individual performances? We carry out this study to assess whether InfoSel can effectively utilize the predictions obtained from various base models, regardless of their individual performance levels. In Figure 3, we observe a minor F1-score difference (0.07%) on the Mini-SDv2 dataset between the InfoSel model ensembled with and without the lowest performing base model (LLaMA). This finding suggests that InfoSel is robust, and not significantly affected by the individual model’s performance. In a more detailed analysis, we observe that InfoSel selects 4% of answers from LLaMA, resulting in an overall gain of +0.28% of the F1-score. This observation highlights the effectiveness of InfoSel, as it can leverage the knowledge contained in the answers provided even by the lowest performing base model to some extent.
Which modality information helps the most for ensembling? In the Table 5, we compare the effect of providing different modality information to InfoSel-MT during ensemble training. Notice that even with just the question and answer (Q+A) information, our model surpasses the performance of the 0-shot PairRanker and LLM-Blender. The setting that yields the lowest accuracy solely utilizes the image (V) as the signal. This can be explained by the fact that a single image often corresponds to multiple questions in VQA datasets, making it challenging for the model to acquire discriminative features. Furthermore, we conclude that a superior performance of our model when utilizing image, question, and answer (V+Q+A) data demonstrates the effectiveness of our model in multimodal setting.
Model | Mini-GQA | Mini-Viz |
(0-shot) PairRanker(Q+A) | 47.69 | 20.74 |
(0-shot) LLM-Blender(Q+A) | 0.0 | 0.0 |
InfoSel-MT(V) | 50.56 | 20.79 |
InfoSel-MT(Q) | 51.11 | 21.21 |
InfoSel-MT(V+Q) | 50.83 | 20.06 |
InfoSel-MT(V+A) | 52.38 | 22.66 |
InfoSel-MT(Q+A) | 54.76 | 22.89 |
InfoSel-MT(V+Q+A) | 55.16 | 23.16 |
4.7 Case Study
Mini-SDv2 | Mini-NQ | |
…Derrick Norman | …in 2005 and the | |
Context: | Lehmer’s list | release of her epon- |
of primes up to | ymous debut album | |
10,006,721… | the following year… | |
@@ 4.7 Case Study
How many primes | When did Taylor | |
Question: | were included in Der- | Swift ’s first |
@@
rick Norman Lehmer’s | album release? | |
list of prime numbers? | ||
LLaMA | unanswerable | 2006 |
Davinci | 10,006,721 | 2006 |
ChatGPT | unanswerable | 2006 |
FT-TT | unanswerable | 2005 |
InfoSel-TT | 10,006,721 | 2006 |
InfoSel∗-TT | unanswerable | 2005 |
Mini-GQA | Mini-Viz | |
Image: | ![]() |
@@
Question: | What appliance is | What is this pro- |
it? | duct? | |
ALBEF | blender | refrigerator |
BLIP | toaster | toilet |
VLMo | microwave | door |
FT-MT | coffee maker | unanswerable |
InfoSel-MT | toaster | toilet |
InfoSel∗-MT | coffee maker | unanswerable |
Table 6 illustrates two insightful cases from the predictions of different models on textual Mini-SDv2 and Mini-NQ QA datasets. The first case showcases the ability of InfoSel-TT to select the right model (Davinci) when the rest of the models is incorrect. However, InfoSel∗-TT selects the wrong answers from the FT-TT model and underperforms InfoSel-TT. The second case illustrates the ability of LLMs to generate correct answer (“2006”) despite the ground truth annotation error (“2005”). This demonstrates the advantage of ensembling highly expressive LLMs instead of relying only on fine-tuning small-size models such as FT-TT.
The first case of Table 7 further indicates that InfoSel captures the only correct answer (“toaster”). The second case demonstrates the ability of InfoSel∗ to recognize task-specific labels (i.e., “unanswerable”) introduced by FT-MT. InfoSel-MT struggles with such labels as they are unfamiliar to the base models. This showcases the benefits of training InfoSel∗ models on datasets containing a high percentage of task-specific labels.
@@ 5 Conclusion
In this paper, we propose InfoSel, a novel lightweight and task-specific ensemble method designed to learn the dynamic selection of the optimal model from a range of distinct black-box base LLMs. We find that using only 110M trainable parameters, our method is able to substantially increase the performance upon the best performing base LLM. Additionally, our analysis reveals that InfoSel remains robust regardless the incorrect predictions of the lowest performing LLM. Our findings also show that our solution is highly data-efficient. Concretely, it requires only a fraction of instances (as few as 10) from the training set to outperform base LLMs. Finally, our experimental results reveal the ability of InfoSel to be adapted to multimodal setting, showing a substantial increase in performance compared to state-of-the-art alternatives.
6 Acknowledgements
This research has been funded by the Vienna Science and Technology Fund (WWTF) [10.47379/VRG19008] “Knowledge-infused Deep Learning for Natural Language Processing”.

### v2

@@ 5.4 Ablation Studies
Is InfoSel robust to the base models’ individual performances? We carry out this study to assess whether InfoSel can effectively utilize the predictions obtained from various base models, regardless of their individual performances. In the first column of each dataset in Figure 5, we observe a minor F1-score difference (0.15%) on the Mini-SDv2 dataset between the InfoSel model ensembled with and without the lowest performing base model (LLaMA). This finding suggests that InfoSel is robust, and not significantly affected by the performance of individual models. In a more detailed analysis, we observe that InfoSel selects 4% of answers from LLaMA on Mini-NQ, resulting in an overall gain of +0.28% of the F1-score. This observation highlights the effectiveness of InfoSel, as it can leverage the knowledge contained in the answers provided even by the lowest performing base model to some extent. However, the results of the second column of each dataset reveal that the ensemble performance decreased significantly when the ChatGPT or Davinci model was removed. This indicates that InfoSel should not be used to ensemble two base models (e.g., LLaMA and ChatGPT) with large performance gaps.
5.5 Case Studies
mini-SDv2 | mini-NQ | |||
Context: | … The building was | … Derrick Norman | Dwight David | … in 2005 and the |
designed by architects | Lehmer’s list | Howard … player | release of her epony- | |
Marek Budzyński and | of primes up to | for the Charlotte | mous debut album | |
Zbigniew Badowski… | 10,006,721 … | Hornets … | the following year … | |
@@ 5.5 Case Studies
Question: | What profession | How many primes were | who did Dwight | when did Taylor |
does Zbigniew | included in Derrick | Howard play | Swift ’s first | |
@@
Marek have? | Norman Lehmer’s list | for last year? | album release? | |
of prime numbers? | ||||
LLaMA-2-70b-chat | architect | unanswerable | Charlotte Hornets | 2006 |
text-davinci-003 | Architect | 10,006,721 | The Houston Rockets | 2006 |
ChatGPT | unanswerable | unanswerable | Washington Wizards | 2006 |
FT-BERT | architects Marek Budzyński | unanswerable | Dwight David Howard | 2005 |
and Zbigniew Badowski | ||||
InfoSel-MT | unanswerable | 10,006,721 | Charlotte Hornets | 2006 |
InfoSel∗-MT | unanswerable | unanswerable | Dwight David Howard | 2005 |
Mini-GQA | Mini-Viz | |||
Image: |
@@
Question: | What appliance is | Is the tall tree on | What kind of food | What is this pro- |
it? | the right? | is in this can? | duct? | |
ALBEF | blender | yes | fruit salad | refrigerator |
BLIP | toaster | yes | vegetable soup | toilet |
VLMo | microwave | yes | fruit | door |
FT-MT | coffee maker | no | soup | unanswerable |
InfoSel-MT | toaster | yes | vegetable soup | toilet |
InfoSel∗-MT | coffee maker | no | vegetable soup | unanswerable |
Table 8 and 9 showcase several interesting cases from the predictions of different models for textual and visual QA tasks. From Table 8, we observe that that InfoSel-TT selects answers from different language models. However, InfoSel∗-TT may select incorrect answers from the overfitted FT-BERT model , leading to underperformance compared to InfoSel-TT in those instances. The final case illustrates a incorrect ground-truth answer provided by the original dataset. Despite this, LLMs generate the correct answer due to their contextual comprehension capabilities. In contrast, FT-BERT, which is limited to classification tasks, can only extract answer tokens from the context and thus fails to provide the correct answer. Therefore, we conclude that leveraging an ensemble of LLMs to harness their robust comprehension abilities offers greater benefits to users than relying on fine-tuning smaller models.
Table 9 demonstrates that both InfoSel-MT and InfoSel∗-MT are capable of identifying the correct answer, even when the prediction of only one of the base models is accurate. The last case demonstrates that InfoSel∗-MT successfully identifies the new label "unanswerable" introduced by FT-MT, a label that InfoSel-MT cannot predict because the base models can not provide. Consequently, incorporating FT-MT into the ensembling training process is crucial when out-of-domain datasets contain a high proportion of new labels.
5.6 Robustness Analysis
We analyze the robustness of the data efficiency property of InfoSel by training InfoSel-TT with 10 different sets of randomly sampled train data. Figure 6 demonstrates the mean and standard deviation of these 10 sets of results on Mini-SDv2 and Mini-NQ. We observe that the deviation decreases as the number of training samples increases. Concretely, when using 1,000 training samples, the deviation is less than 0.5 (0.29 for Mini-SDv2, 0.40 for Mini-NQ). Remarkably, the mean F1-score (61.12) achieved with just 10 samples from Mini-SDv2 surpasses that of the best base model. Furthermore, the mean F1-score (71.60) obtained with 100 samples from Mini-NQ exceeds the F1-score of the best base model, despite using fewer samples than the 500 samples reported in the main result figure (Figure 3).
5.7 Discussion and Limitations
InfoSel offers an effective approach to enhancing out-domain black-box model performance and addressing answer selection. However, it is important to acknowledge certain limitations associated with its application, which are outlined in the following paragraphs.
Dependency on Annotated Data: InfoSel, similar to numerous machine learning techniques, relies on a small amount of annotated training and development data tailored to a specific domain. While this requirement is relatively modest, and the efficiency of InfoSel in data utilization has been demonstrated through experiments, it may still pose a limitation in scenarios where obtaining such data is challenging or costly.
Limited Applicability to Open-Ended Text Generation: the primary strength of InfoSellies in its ability to select the best answer from a set of base models, making it particularly valuable in question-answering scenarios. However, for more open-ended text-generation tasks, where it may be beneficial to combine multiple answers, InfoSel’s single-answer selection mechanism may not be the ideal choice, and future research directions may include approaches for combining several long-form answers.
Transparency and Explainability: InfoSel, similar to other machine learning models that derive answers from black-box systems, may also function as a "black box". This implies that its decision-making process may not be easily interpretable or understandable to end-users. By integrating InfoSel with explainability techniques, users may gain a more transparent view of how the model arrives at its selections.
@@ 6 Conclusion
In this paper, we propose InfoSel, a novel lightweight and task-specific ensemble method designed to learn the dynamic selection of the optimal answer from a range of distinct black-box base models. InfoSel is able to handle multimodal data and we demonstrate its effectiveness on textual and visual question answering tasks.
Our findings also show that InfoSelis highly data-efficient and, using only a fraction of the trainable parameters relative to the base models, our method can substantially increase the performance over the best-performing base LLMs and VQA models. Concretely, it requires only a fraction of instances from the training set to outperform base LLMs and previous ensembling methods. An in-depth exploration of the importance of different input information for ensembling shows that fused (rather than concatenated) information of diverse types (image, question, answer) achieved the best performance. An ablation and robustness analysis reveals that InfoSel is robust to incorrect predictions of low-performing LLMs and to the randomness of training data selection.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The data used in this paper is publicly accessible.
Acknowledgements
This research has been funded by the Vienna Science and Technology Fund (WWTF) [10.47379/VRG19008] “Knowledge-infused Deep Learning for Natural Language Processing”. Co-funded by the European Union (grant Agreement no. 101146515).999Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.